Circular No. 7787

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1999 TC_{36}

C. A. Trujillo and M. E. Brown, California Institute of Technology, report the discovery of a companion to the plutino 1999 TC_{36} (MPEC 1999-Y19, 2000-O37; MPO 20742). Observations were collected on two dates, the first spanning ∼ 45 min centered on 2001 Dec. 8.906 UT and the second on Dec. 9.307, as part of an ongoing survey for transneptunian binaries undertaken with the Hubble Space Telescope STIS instrument. Separations of 0′′.3651 ± 0′′.0011 in p.a. 2^{o}.0 ± 0^{o}.1 and 0′′.3698 ± 0′′.0007 in p.a. 3^{o}.7 ± 0^{o}.1 were measured for the two dates, respectively. At a geocentric distance of 31.16 AU, this separation corresponds to at least 8000 km between components. The pair exhibit a flux difference of 2.21 ± 0.01 magnitudes in the unfiltered images. Recovery observations are necessary for companion orbit determination; however, due to the close proximity and large flux difference between the primary and companion, uncorrected ground-based recovery is likely to be difficult.

This report provides an opportunity to correct a misconception on IAUC 7749. In referring to IAUC 7733, it was suggested that 2001 QT_{297} is a plutino, whereas it is more likely to be a cubewano.

SUPERNOVAE 2001is AND 2001it T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001it (cf. IAUC 7782), obtained by M. Calkins on Jan. 6.55 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. Weak P-Cyg lines of hydrogen are apparent on a relatively blue continuum.

S. Benetti, G. Altavilla, A. Pastorello, M. Turatto, S. Desidera, and E. Giro, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write: “A spectrum (range 400–780 nm, resolution 2.5 nm) of SN 2001is (cf. IAUC 7782), obtained on Jan. 8.86 UT with the Asiago 1.82-m telescope (+ AFOSC), shows it to be a type-Ib supernova, 2–3 weeks after maximum. The spectrum closely resemble those of the type-Ib SN 1999dn, which has a residual hydrogen contamination in the spectra (Branch et al. 2002, Ap.J., in press). We also observed SN 2001it, whose spectrum is that of a type-II supernova around maximum; Hα shows a broad emission feature, without absorption, from which an expansion velocity (FWHM) of ∼ 10 500 km/s is derived.”

2002 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green